# Agent-Oriented Planning in Multi-Agent Systems

### v1

@@ 3.3 Sub-Task Modifications
We identify two intrinsic reasons for such a situation in agent-oriented planning. First, the description of the sub-task may lack critical information necessary for solving the problem, such as the interpretation of contextually elements like pronouns or other ambiguous terms. In such cases, while the agent possesses the capability to resolve the sub-task, the missing information can hinder it to provide a satisfactory response. Second, the sub-task $q_{i}$ might be too complex, therefore a single agent can only address parts of this sub-task based on its expertise and tools. In this subsection, we design mechanisms to distinguish between these two cases and propose corresponding solutions.
To be specific, for each agent, we propose to construct a set of representative works that record the tasks it has completely resolved. We would initially bypass the construction and update for these representative works (refer to Section 3.5 for more details) and focus on how to utilize them to make sub-tasks modifications here. Given a sub-task $q_{i}$, we calculate its similarity with a representative work $q_{t}$ following $cos(\mathcal{E}(q_{i}),\mathcal{E}(q_{t}))$, where $\mathcal{E}$ is the embedding part in $\mathcal{M}$ and $cos(u,v)=\frac{u\cdot v}{\left\|u\right\|_{2}\left\|v\right\|_{2}}$ calculates the cosine similarity. We define the similarity as follows:
@@
On the other hand, if the calculated similarity does not meet the threshold, we regarded $q_{i}$ as too complex for any single agent in the system to solve (corresponding to the second reason above). As a result, the meta-agent $\mathcal{P}$ is requested to perform plan-in-detail for further decomposing the sub-task into simple ones. To avoid creating redundant sub-tasks, we provide $\mathcal{P}$ with all the sub-tasks and explicit instruction on avoiding overlapping sub-asks. The prompt used for sub-task modification can be found in Appendix A.3 and Appendix A.4.
@@ 3.4 Detector for Completeness and Non-Redundancy
Following the idea in fast decomposition and allocation, we attempt to enhance the proposed framework by incorporating detailed instructions to satisfy the principles of completeness and non-redundancy, such as “make sure all the important information of the original task such as nouns or numbers are included in the sub-tasks”. However, empirical observations reveal that these instructions do not mitigate these issues. In fact, more than 15% of user queries still exhibit such issues during decomposition.
To tackle these issues, we involve a detector, implemented with providing a role-play prompt to LLMs, to evaluate both the completeness and non-redundancy of the intermediate results provided by the meta-agent. For the principle of completeness, the detector extracts all key elements and requirements from the original task, and then matches these elements against each sub-task, determining whether the sub-tasks collectively address all essential aspects of the original task. Regarding the principle of non-redundancy, the detector checks for identical information and requirements across sub-tasks. The detector flags the redundancy if two sub-tasks include overlapping content.
When the provided results do not satisfy the principles of completeness or non-redundancy, the detector identifies the missing key information or redundant content, and offers recommendations for refining the plan, such as suggestions for supplementing missing details or removing overlapping sub-tasks.
The prompt used for the detector is provided in Appendix A.5. From the experimental results shown in Section 2, we observe that the detector effectively mitigates the issues of task decomposition incompleteness and redundancy, leading to an effective and efficient agent-oriented planning process.
@@ 3.5 Feedback Loop
An automatic feedback loop is integrated into the proposed agent-oriented planning framework, promoting ongoing enhancement of the meta-agent.
As introduced in Section 3.2, the reward model can identify certain sub-tasks as being resolvable by specific agents. These identified sub-tasks are then collected to form the representative works of the corresponding agents. The training data for the reward model, which includes the sub-tasks and the ground truth scores of responses provided by the agents, is utilized to initialize these representative works.
Note that the representative works are continuously updated. When a user query is completely resolved, the sub-tasks decomposed from that user query can be added to the representative works of the agents who are selected to provide response to these sub-tasks. To prevent redundancy in the representative works, we implement a similarity threshold. This threshold ensures that only new and sufficiently distinct sub-tasks are incorporated into an agent’s representative works, maintaining the diversity and relevance of the tasks that each agent has previously resolved.
@@ 4 Experiments
In this section, we provide empirical comparisons between the proposed agent-oriented planning framework and the existing studies.
@@ 4.1 Settings
We conduct experiments based on a numerical reasoning dataset (Kim et al., 2024), which necessitates the collaboration of multiple agents in resolving the queries. For example, a query can be “If Sarah wants to buy one BMW X5 and one Tesla Model 3, how much more would she need to pay to buy the BMW X5 compared to the Tesla Model 3?” To resolve this query, we first need to search for the prices of the BMW X5 and Tesla Model 3, and then calculate the price difference between them. Following previous study (Kim et al., 2024), we construct a training dataset consisting of 1,440 queries, along with an annotated test dataset consisting of 292 queries. Note that the topics in the training dataset and test dataset may vary slightly.
For quantification comparisons, we provide instructions to GPT-4o (OpenAI, 2024b) for assisting in judging whether the execution results align with the ground truth, and calculate the accuracy as the evaluation metrics. The prompt used for evaluation can be found in Appendix D.
@@
The meta-agent is tasked with decomposing user queries into sub-tasks and assigning the most suitable agents to execute those sub-tasks. We use GPT-4o as the employed LLM for all agents, and the used prompt are provided in Appendix B.
@@
We adopt all-MiniLM-L6-v2 (Wang et al., 2020) as the embedding layers of the reward model, which maps text sentences or paragraphs into a 384-dimensional dense vector space. The embeddings for both the sub-task and the agent’s description are computed separately, and are concatenate together to form a 768-dimensional dense vector. Following the embedding model, we add a three-layer multilayer perceptron (MLP) with output dimensions of 256,64, and 1, respectively. We freeze the embedding layers and only finetune the parameters of the MLP, making the training process efficiently. The batch size are set to 32, the learning rate is 1e-3. We train the reward model for 50 epochs on one Tesla V100-SXM2-32GB GPU.
@@
For the single-agent systems, we utilize GPT-4o as a baseline method, providing user queries directly without any additional prompt engineering. Besides, we provide GPT-4o with Chain-of-Thought (CoT) (Wei et al., 2022) prompt and Zero-Shot CoT (Kojima et al., 2022) prompt to guide LLMs to perform reasoning processes, resulting in two baselines denoted as CoT and Zero-Shot CoT. Taking a further step, we instruct GPT-4o to serve as a meta-agent responsible for performing fast task decomposition and allocations. We implement two baselines, denoted as Meta-Agent and Meta-Agent: Traversal, respectively. Meta-Agent denotes that the LLM selects one agent for each sub-task, while Meta-Agent: Traversal denotes that GPT-4o iteratively queries all agents for each sub-task and selects the best response.
Regarding multi-agent systems, we compare the proposed framework with REACT (Yao et al., 2022) and HUSKY (Kim et al., 2024), which are representative methods involving task decomposition and sub-task execution.
Table 1: Comparisons between the proposed agent-oriented planning framework and baselines.
@@ 4.2 Comparisons and Analysis
The comparisons between the proposed agent-oriented planning framework and the baseline methods are shown in Table 1. We adopt accuracy as the metric for comparing effectiveness, and use prompt tokens, completion token, and the execution time for comparing efficiency.
Overall, the experimental results demonstrate that the proposed agent-oriented planning framework achieves notable improvements compared to all baseline methods. To be specific, when comparing with single-agent systems including GPT-4o, CoT, and Zero-Shot CoT, the proposed agent-oriented planning framework outperforms these systems by 10.4%, 8.1%, and 11.5% in terms of accuracy, respectively. These improvements can be attributed to the collaboration among multiple diverse agents and the effective scheduling provided by the meta-agent. It is not surprising to observe that the proposed framework incurs higher resource costs and longer inference times compared to single-agent systems. These additional costs during the inference phase, which includes task decomposition, allocation, and modifications carried out by the meta-agent, are affordable and can be worthwhile as long as they bring significant improvements in accuracy and stability when applying real-world applications.
Compared to systems that also involve a meta-agent for task decomposition and allocation, we can observe from the table that the proposed method achieves at least a 4.1% improvement in terms of accuracy while maintaining the same level of computation costs and inference times. The results of Meta-Agent and Meta-Agent: Iteration indicate that simply instructing GPT-4o to perform task decomposition and allocation does not always yield satisfactory responses, even though the capabilities of LLM are recognized as powerful. Existing studies in task composition and allocation fail to consider the abilities and characteristics of the agents beyond their descriptions, and lack mechanisms for modifications and feedback within the multi-agent system, which can be hindered by the challenges as summarized in Section 2. The proposed framework is built upon these identified design principles and incorporates novel mechanisms for automatically evaluating intermediate results, allowing timely modifications and revisions, which guides and leverages the strengths of LLMs within the multi-agent system for effectively resolving user queries and enhancing overall system performance.
@@ 4.3 Further Discussions
We conducted an ablation study to confirm the contributions of different components in the proposed framework. Specifically, we disable the detector, the reward model, and the representatives of agents in separate experiments, with results reported in Table 2. From these results, we can observe that the detector has a significant impact on execution accuracy, as indicated by a notable increase in the incompleteness rate during the meta-agent’s task decomposition phase. Both the reward model and the representative works are necessary in ensuring the solvability of tasks and in selecting the most suitable agents for each sub-task. Overall, these results demonstrate that all three components are indispensable, working together to ensure the feasibility of task decomposition and allocation, leading to satisfactory responses to the original user queries.
@@
Impact of Scorer and Reward Model
The training dataset is constructed based on annotations provided by the scorer, which in the previous experiments is implemented using LLMs. In this section, we shift to using human annotators to manually provide scores for the responses, evaluating whether a reward model trained on this manually labeled dataset would further enhance performance. This approach is denoted as Manual Scoring. More detailed information about the manual scoring process can be found in Appendix C. Besides, based on the manually scored training dataset, we investigate the setting of updating the embedding layers of the reward model, denoted as Full Parameter Tuning.
The experimental results shown in Table 3 indicate that both manual scoring and full parameter tuning of the reward model lead to improved performance. These results suggest the potential for further improving the proposed agent-oriented planning framework by providing high-quality datasets for training robust reward models.
@@
5 Related Work
The field of artificial general intelligence has witnessed a transformative development with the advent of large language models (LLMs), such as GPT-3 (Brown, 2020), GPT-4 (Achiam et al., 2023), PaLM (Chowdhery et al., 2023), LLaMa (Touvron et al., 2023) and so on (Devlin et al., 2018; Zhang et al., 2022; Chowdhery et al., 2023; Le Scao et al., 2023). LLMs have demonstrated remarkable capabilities in understanding and generating human language (Zhao et al., 2023), and solving tasks such as mathematical problems and commonsense reasoning (Imani et al., 2023; Ahn et al., 2024). To further enhance the ability of LLMs in handling complex problems, some research focuses on improving the internal reasoning capabilities of LLMs, typically involves decomposing complex questions into sequential intermediate steps before generating the final answer, as exemplified by Chain-of-Thought (CoT) prompting (Wei et al., 2022) and its variants (Kojima et al., 2022; Wang et al., 2022; Zhou et al., 2022). Some methods opt to generate a dozen possible paths to comprise the candidate plan set (Yao et al., 2024; Besta et al., 2024; Zhao et al., 2024; Hao et al., 2023) and select the optimal plan with strategies like heuristic search algorithms (Xiao & Wang, 2023). In environments with complex constraints, such as generating permissible actions, several methods combine LLMs with external planners such as symbolic planners (Liu et al., 2023; Dagan et al., 2023; Guan et al., 2023) and neural planners (Yao et al., 2020; Lin et al., 2024).
Despite these advancements, LLMs still face challenges in addressing tasks that require extensive knowledge, such as avoiding hallucinations, integrating diverse domain knowledge, and executing long-horizon planning (Qin et al., 2023a; Bubeck et al., 2023; Maynez et al., 2020). To address these issues, recent work has focused on multi-agent systems, which leverage the collective intelligence and specialized skills of multiple LLM-empowered agents to solve tasks. Compared with reinforcement learning, LLM-empowered agents have more comprehensive internal world knowledge, and can provide natural language interfaces to interact with humans, which is more flexible and explainable (Wang et al., 2024). Serving as the main controller or “brain” that manages the flow of operations, the meta-agent has gained the most significant attention. The meta-agent is responsible for understanding the overall task, decomposing it into smaller ones, and coordinating the sub-tasks to the agents. There are mainly two technical approaches in task decomposition: interleaved decomposition and decomposition-first. In the first kind of method, React (Yao et al., 2022) generates reasoning and action in an interleaved manner, where reasoning helps the meta-agent update action plans and action enables interaction with agents. Visual ChatGPT (Wu et al., 2023a) meticulously designs a series of prompts to help inject the visual information into ChatGPT, designed to solve complex visual questions with diverse visual foundation models step-by-step. HUSKY (Kim et al., 2024) trains an action model to iterate between two stages: generating the next action to take and executing the action using expert models and updating the current solution state.
Regarding complex tasks, excessively long trajectories might bring hallucinations. Given that, HuggingGPT (Shen et al., 2024) carefully designs prompts to instruct ChatGPT to decompose tasks, selecting models in HuggingFace and summarizing the response. AutoAgents (Chen et al., 2023) dynamically generates multiple required agents based on task content and plans solutions for the current task based on the generated expert agents. Chameleon (Lu et al., 2024) prompts the meta-agent with tool descriptions and usage examples to infer a program composed of a sequence of tools to execute in order to generate the final response.
Although remarkable progress has been made, plans generated solely through prompt-based methods often fall short of meeting the principles of solvability, completeness, and non-redundancy—issues that have largely been overlooked in prior work. Besides, the meta-agent’s ability to make suitable schedules can be hindered when relying only on agents’ descriptions without considering their actual capabilities or past performance. To address these issues, we propose a novel framework to perform task decomposition and allocation satisfying the aforementioned three principles.
@@ 6 Conclusion
In this study, we propose a novel agent-oriented planning framework for multi-agent systems, following three critical design principles to ensure that the meta-agent can effectively decompose the user query into several sub-tasks for producing satisfactory responses. The proposed framework utilizes a fast decomposition and allocation process, which relies on the ability of LLMs to generate an intermediate schedule efficiently. After that, a reward model and the representative work mechanism are employed to evaluate these intermediate results, routing three different paths for executing the sub-task or making necessary modifications to align the sub-task with agents, such as replan the sub-task, plan in detail, and re-describe. Extensive experiments demonstrate that the proposed agent-oriented framework achieves significant improvements over both the existing single-agent and multi-agent baseline methods. We provide discussions on the contributions of different components in the framework and the potential for improvements in agent-oriented planning. We will release the source code to promote further research in the community.
Appendix A Plan Generation and Modification
A.1 Meta-agent Prompt

### v2

@@ 4.3 Sub-Task Modifications
We identify two intrinsic reasons for such a situation in agent-oriented planning. First, the description of the sub-task may lack critical information necessary for solving the problem, such as the interpretation of contextual elements like pronouns or other ambiguous terms. In such cases, while the agent possesses the capability to resolve the sub-task, the missing information can hinder it from providing a satisfactory response. Second, the sub-task $q_{i}$ might be too complex, therefore a single agent can only address parts of this sub-task based on its expertise and tools. In this subsection, we design mechanisms to distinguish between these two cases and propose corresponding solutions.
To be more specific, for each agent, we propose to construct a set of representative works that record the sub-tasks it has completely resolved. We would initially bypass the construction and update for these representative works (refer to Section 4.5 for more details) and focus on how to utilize them to make sub-tasks modifications here. Given a sub-task $q_{i}$, we calculate its similarity with a representative work $q_{t}$ following $cos(\mathcal{E}(q_{i}),\mathcal{E}(q_{t}))$, where $\mathcal{E}$ is the embedding part in $\mathcal{M}$ and $cos(u,v)=\frac{u\cdot v}{\left\|u\right\|_{2}\left\|v\right\|_{2}}$ calculates the cosine similarity. We define the similarity as follows:
@@
On the other hand, when the calculated similarity does not meet the threshold, we regard $q_{i}$ as too complex for any single agent in the system to solve (corresponding to the second reason above). As a result, the meta-agent $\mathcal{P}$ is requested to perform plan-in-detail for further decomposing the sub-task into simple ones. To avoid creating redundant sub-tasks, we provide $\mathcal{P}$ with all the sub-tasks and explicit instruction on avoiding overlapping sub-asks. The prompt used for sub-task modification can be found in Appendix A.3 and Appendix A.4.
@@ 4.4 Detector for Completeness and Non-Redundancy
Following the idea in fast decomposition and allocation, we attempt to incorporate detailed instructions to the meta-agent to satisfy the principles of completeness and non-redundancy. However, empirical observations reveal that these instructions may not always work well. In fact, more than 15% of user queries still exhibit such issues during decomposition in our initial experiments.
To tackle these issues, we involve a detector, implemented by providing a role-play prompt to LLMs, to evaluate both the completeness and non-redundancy of the intermediate results provided by the meta-agent. For the principle of completeness, the detector extracts all key elements and requirements from the original query and then matches these elements against each sub-task, determining whether the sub-tasks collectively address all essential aspects of the original task. To further ensure that each sub-task can be successfully executed, the detector is also required to determine whether there are any additional dependencies needed beyond those identified during decomposition (e.g., unforeseen dependencies). If such dependencies exist, the execution results of these dependent sub-tasks should also be provided as inputs.
Regarding non-redundancy, the detector is responsible for analyzing whether any two subtasks contain identical information or try to solve the same problem, and whether there are unnecessary subtasks that do not contribute to the user query’s solution.
When the provided results do not satisfy the principles of completeness or non-redundancy, the detector identifies the missing key information or redundant content and offers recommendations for refining the plan, such as suggestions for supplementing missing details or removing overlapping sub-tasks. The prompt used for the detector is provided in Appendix A.5. From the experimental results shown in Table 2, we observe that the detector effectively mitigates the issues of incompleteness and redundancy, leading to an effective and efficient agent-oriented planning process.
@@ 4.5 Feedback Loop
An automatic feedback loop is integrated into AOP, promoting the ongoing enhancement of the problem-solving process. As introduced in Section 4.2, the reward model can identify certain sub-tasks as being resolvable by specific agents. These identified sub-tasks are then collected to form the representative works of the corresponding agents. The training data for the reward model, which includes the sub-tasks and the ground truth scores of responses provided by the agents, is utilized to initialize these representative works.
Note that the representative works are continuously updated. When a user query is completely resolved, the sub-tasks decomposed from that user query can be added to the representative works of the agents who are selected to provide responses to these sub-tasks. To prevent redundancy in the representative works, we implement a similarity threshold. This threshold ensures that only new and sufficiently distinct sub-tasks are incorporated into an agent’s representative works, maintaining the diversity and relevance of the tasks that each agent has previously resolved.
@@ 5 Experiments
In this section, we provide empirical comparisons between AOP and the existing studies.
@@ 5.1 Settings
We conduct experiments based on a numerical reasoning dataset (Kim et al., 2024), which necessitates the collaboration of multiple agents in resolving the queries. For example, a query can be “If Sarah wants to buy one BMW X5 and one Tesla Model 3, how much more would she need to pay to buy the BMW X5 compared to the Tesla Model 3?” To resolve this query, we first need to search for the prices of the BMW X5 and Tesla Model 3, and then calculate the price difference between them. Following previous study (Kim et al., 2024), we adopt Husky-QA, which consists of 1,440 queries in the training data and 292 queries in the test data. Besides, we also provide more experimental results on the decontextualized versions of a subset of DROP (Dua et al., 2019) and IIRC (Ferguson et al., 2020) in Appendix D.1.
For quantification comparisons, we provide instructions to GPT-4o (OpenAI, 2024) for assisting in judging whether the execution results align with the ground truth and calculate accuracy as the evaluation metric. The prompt used for evaluation can be found in Appendix A.6.
@@
The meta-agent is tasked with decomposing user queries into sub-tasks and assigning the most suitable agents to execute those sub-tasks. We use GPT-4o as the employed LLM for all agents, and the prompts used are provided in Appendix A.7. Besides, an investigation on the effectiveness of involving expert agents powered by task-specified LLMs, and the method for efficiently handling the coexistence of similar agents can be found in Appendix E.
@@
We adopt all-MiniLM-L6-v2 (Wang et al., 2020) as the embedding layers of the reward model, which maps text sentences or paragraphs into a 384-dimensional dense vector space. The embeddings for both the sub-task and the agent’s description are computed separately, and then are concatenated together to form a 768-dimensional dense vector. Following the embedding model, we add a three-layer multilayer perceptron (MLP) with output dimensions of 256,64, and 1, respectively. We freeze the embedding layers and only fine-tune the parameters of the MLP, making the training process efficient. The batch size is set to 32, and the learning rate is 1e-3. We train the reward model for 50 epochs on one Tesla V100-SXM2-32GB GPU.
@@
For the single-agent systems, we utilize GPT-4o as a baseline method, providing user queries directly without any additional prompt engineering. Besides, we provide GPT-4o with Chain-of-Thought (CoT) (Wei et al., 2022) prompt and Zero-Shot CoT (Kojima et al., 2022) prompt to guide the LLM to perform reasoning processes, resulting in two baselines denoted as CoT and Zero-Shot CoT.
Regarding the multi-agent systems, we instruct GPT-4o to serve as a meta-agent responsible for performing fast task decomposition and allocation. We implement two baselines, denoted as Meta-Agent and Meta-Agent: Traversal, respectively. Meta-Agent denotes that GPT-4o selects one agent for each sub-task, while Meta-Agent: Traversal denotes that GPT-4o iteratively queries all agents for each sub-task and selects the best response. Besides, we compare AOP with REACT (Yao et al., 2023) and HUSKY (Kim et al., 2024), which are representative methods involving task decomposition and sub-task execution.
Table 1: Comparisons between AOP and baselines.
@@ 5.2 Comparisons and Analysis
The comparisons between AOP and the baseline methods are shown in Table 1. We adopt accuracy as the metric for comparing effectiveness, and use prompt/completion tokens (refer to the total cost on the whole test dataset) and the execution time for comparing efficiency.
Overall, the experimental results demonstrate that AOP achieves notable improvements compared to all baseline methods. To be specific, when comparing with single-agent systems including GPT-4o, CoT, and Zero-Shot CoT, AOP outperforms these systems by 10.4%, 8.1%, and 11.5% in terms of accuracy, respectively. These improvements can be attributed to the collaboration among multiple diverse agents and the effective scheduling provided by the meta-agent.
It is not surprising to observe that the cost of AOP is significantly higher than that of single-agent systems, and is at the same level compared to other multi-agent systems. These additional costs during the inference phase, which include task decomposition, allocation, and modifications carried out by the meta-agent, are affordable and can be worthwhile as long as they bring significant improvements in accuracy and stability when applied to real-world applications. Such an exploration aligns with a recent study (OpenAI, 2024) in inference time computing, aimed at efficiently utilizing more tokens to resolve challenging tasks.
Compared to systems that also involve a meta-agent for task decomposition and allocation, we can observe from the table that AOP achieves at least a 4.1% improvement in terms of accuracy while maintaining the same level of computation costs and inference times. The results of Meta-Agent and Meta-Agent: Iteration indicate that simply instructing GPT-4o to perform task decomposition and allocation does not always yield satisfactory responses, even though the capabilities of LLMs are recognized as powerful. Existing studies in task decomposition and allocation fail to consider the abilities and characteristics of the agents beyond their descriptions, and lack mechanisms for modifications and feedback within the multi-agent system, which can be hindered by the challenges as summarized in Section 3. AOP is built upon these identified design principles and incorporates novel mechanisms for automatically evaluating intermediate results, allowing timely modifications and revisions, which guide and leverage the strengths of agents in the system to effectively resolve user queries and enhance the overall system performance.
@@ 5.3 Further Discussions
We conducted an ablation study to confirm the contributions of different components in AOP. Specifically, we disable the detector, the reward model, and the representative works of agents in separate experiments, with results reported in Table 2. From these results, we can observe that the detector has a significant impact on execution accuracy, as indicated by a notable increase in the incompleteness rate during the meta-agent’s task decomposition phase. Both the reward model and the representative works are necessary for ensuring the solvability of sub-tasks and for selecting the most suitable agents for each sub-task. Overall, these results demonstrate that all three components are indispensable, working together to ensure the feasibility of task decomposition and allocation, leading to satisfactory responses to the original user queries.
@@
Impact of the Scorer and Reward Model
The training dataset is constructed based on annotations provided by the scorer, which in previous experiments was implemented using LLMs. In this section, we shift to using human annotators to manually provide scores for the responses, evaluating whether a reward model trained on this manually labeled dataset would further enhance performance. This approach is denoted as Manual Scoring. More detailed information about the manual scoring process can be found in Appendix C. Besides, based on the manually scored training dataset, we investigate the setting of updating the embedding layers of the reward model, denoted as Full Parameter Tuning.
The experimental results shown in Table 3 indicate that both manual scoring and full parameter tuning lead to improved performance. These results suggest the potential for further improving AOP by providing high-quality datasets for robust training. Although using a human-expert-based scorer can lead to further improvements, we predominantly adopt a model-based scorer in the above experiments, as it is more cost-effective and generalizable. Besides, we provide empirical studies on the generalization capability of the reward model in Appendix D.2, showing that the reward model trained on one dataset demonstrates good generalization to other datasets.
@@ 6 Conclusion
In this study, we propose AOP, a novel agent-oriented planning framework for multi-agent systems, following three critical design principles to ensure that the meta-agent can effectively decompose the user query into several sub-tasks for producing satisfactory responses. AOP utilizes a fast decomposition and allocation process, which relies on the ability of LLMs to generate an intermediate schedule efficiently. After that, a reward model and the representative works mechanism are employed to evaluate these intermediate results, resulting in executing the sub-task or making necessary modifications to align the sub-task with agents, such as replan the sub-task, plan in detail, and re-describe. Extensive experiments demonstrate that AOP achieves significant improvements over both the existing single-agent and multi-agent baseline methods. We provide discussions on the contributions of different components in AOP and the potential for improvements in agent-oriented planning. We hope AOP can contribute to advancing research in this field, inspiring further exploration and innovation in practical applications.
Appendix A Prompts Used in AOP
A.1 Meta-agent